OpenStudy (anonymous):

Guys, the wolframalpha answer doesn't look like mine, need help to find where is my mistake for this integration: ( 1- sqrt(x) / 1 + sqrt(x) )

OpenStudy (anonymous):

\[2\int\limits_{}^{} -1/u + 1/( (u-1)(u)) du\] sorry this should my second post

OpenStudy (anonymous):

\[2∫−2/u+1/(u+1)du\]

OpenStudy (anonymous):

\[-4\ln(1+\sqrt(x)) + 2\ln (\sqrt(x))\]

OpenStudy (anonymous):

Refer to the WolframAlpha attachment.